# Retrieval Augmented Diffusion Model for Structure-informed Antibody Design and Optimization

### v1

@@ 1 Introduction
Antibodies, essential Y-shaped proteins in the immune system, are pivotal for recognizing and neutralizing specific pathogens known as antigens. This specificity primarily arises from the Complementarity Determining Regions (CDRs), which are crucial for binding affinity to antigens (Jones et al., 1986; Ewert et al., 2004; Xu & Davis, 2000; Akbar et al., 2021). The design of effective CDRs is therefore central to developing potent therapeutic antibodies, a dominant class of protein therapeutics. However, the development of these antibodies typically relies on labor-intensive experimental methods such as animal immunization or screening extensive antibody libraries, often failing to produce antibodies that target therapeutically relevant epitopes effectively. Thus, the ability to generate new antibodies with pre-defined biochemical properties in silico carries the promise of speeding up the drug design process.
Computational efforts in antibody design have traditionally involved grafting residues onto existing structures (Sormanni et al., 2015), sampling alternative native CDR loops to enhance affinities(Aguilar Rangel et al., 2022), and using tools like Rosetta for sequence design improvements in interacting regions (Adolf-Bryfogle et al., 2018). Many recent studies have focused on applying deep generative models to design antibodies (Luo et al., 2022; Martinkus et al., 2024; Zhu et al., 2024). They take advantage of geometric learning and generative models to capture the higher-order interactions among residues directly from the data. These innovations provide more efficient methods to search sequence and structure spaces.
Albeit powerful, current generative models struggle to design antibodies that adhere to structural constraints and exhibit desired biological properties. This challenge primarily arises from a lack of diversity in the available training data. Predominantly, research efforts have relied on the SAbDab database (Dunbar et al., 2014), which comprises fewer than ten thousand antigen-antibody complex structures. The limited scope of this dataset restricts the models’ ability to capture comprehensive high-order interaction information between antigen-antibody residues, thereby increasing the risk of overfitting. Moreover, most existing methodologies attempt to design antibody sequences de novo, without the benefit of template-based guidance. This approach inherently demands a greater volume of data and extensive training or fine-tuning on specific datasets to achieve efficacy in practical applications.
@@
To this end, we introduce the Retrieval-Augmented Diffusion Antibody design model (RADAb), a novel semi-parametric antibody design framework. To fully exploit the protein structure space, we first compiled a database of CDR-like fragments from the non-redundant Protein Data Bank (PDB) (Berman et al., 2000). These CDR-like fragments are linear functional motifs structurally compatible with an antibody CDR loop, found in any protein geometry within the PDB. For a given antibody to be improved, we perform a structural retrieval to obtain motifs with structures similar to the desired CDR framework. As protein sequences capable of folding into similar structures often share homology and consensus, we hypothesize that these retrieved motifs, enriched with evolutionary information, can enhance the model’s generalization.
Unlike traditional rational design methods that optimize by grafting a single CDR fragment, we propose to use a set of structural homologous CDR-like motifs together with the desired backbone for iterative sequence optimization (Figure 1). Our major contributions follow: (1) We propose a first-of-its-kind retrieval-augmented generative framework for rational antibody design. It uses a set of functional CDR-like fragments that satisfy the desired backbone structures and properties to guide generation toward satisfying all the required properties. (2) A novel retrieval mechanism is introduced for integrating these exemplar motifs with the input backbone through a novel dual-branch denoising module, utilizing both structural and evolutionary information. Additionally, we present a coupled conditional diffusion module that iteratively refines the evolution process by incorporating global and local conditions. This allows the model to incorporate more functional information than traditional antibody inverse folding models. (3) Empirical experiments demonstrate that our method improves the state-of-the-art methods in multiple antibody inverse folding tasks, e.g., an 8.08% AAR gains in long CDRH3 inverse folding task and an average of 7 cal/mol absolute $\Delta$$\Delta$G improvements in functionality optimization task, offering a fresh perspective on biomolecular generative models.
@@ 2 Related work
Antibody Design Computational antibody design primarily follows two paths: conventional energy function optimization methods and machine learning approaches. Early antibody design methods were often limited to sequence similarity and energy function optimization (Lapidoth et al., 2015; Adolf-Bryfogle et al., 2018). Recent success of machine learning approaches mainly falls into two directions: antibody sequence design and antigen-specific antibody sequence-structure co-design. The methods used for antibody sequence design mainly include language-based models (Ruffolo et al., 2021; Olsen et al., 2022; Wang et al., 2023a) and inverse folding models (Dreyer et al., 2023; Høie et al., 2024). The other line focuses on antibody sequence-structure co-design mainly taking antibody-antigen complex as a graph, then using graph networks to extract features and predict the coordinates and residue type of antibody CDR (Jin et al., 2021; Kong et al., 2022; 2023; Lin et al., 2024; Luo et al., 2022; Zhu et al., 2024; Martinkus et al., 2024). Despite powerful, these works typically create antibodies from scratch without template constraints, leading to model optimization challenges and poor generative capability. Instead, our method leverages the power of templates from a structure-informed perspective.
@@
Diffusion generative models Diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020) are a class of generative models that have achieved impressive progress on a lot of generation tasks. Denoising diffusion probabilistic models (DDPMs) are a branch of diffusion models, which contain two Markov processes. The forward process perturbs the data into pure noise, and then learns to generate data by reversing the forward Markov process. Because of the diffusion model’s flexibility and controllability, numerous works are focusing on employing retrieval-augmented methods to complement the diffusion framework for text-to-image generation (Sheynin et al., 2023), image generation (Blattmann et al., 2022), human motion generation (Zhang et al., 2023a) and small molecule generation (Wang et al., 2023b; Huang et al., 2024).
@@
Retrieval augmented generative models Retrieval augmented generation technique was first proposed in the field of natural language processing to enhance the language models by introducing an additional database (Lewis et al., 2020; Guu et al., 2020), prompting the language models to generate more realistic and diverse results. Subsequently, retrieval augmented generation (RAG) has conducted diverse explorations in large fields, including natural language processing (Zhang et al., 2023b; Gao et al., 2023; Xu et al., 2024; Yoran et al., 2024; Caffagni et al., 2024) and computer vision (Long et al., 2022; Blattmann et al., 2022; Hu et al., 2023; Rao et al., 2023). Although there has been some retrieval work in the field of protein design and discovery (Zhou & Grigoryan, 2015; Aguilar Rangel et al., 2022), to the best of our knowledge, this is the first-of-its-kind retrieval-based generative framework for antibody design.
@@ 3.1 Notations
Antibody consists of two heavy chains and two light chains. Each chain’s tip has a complementary site that specifically binds to a unique epitope on the antigen. This site includes six complementary-determining regions (CDRs): CDR-H1, CDR-H2, and CDR-H3 on the heavy chain, and CDR-L1, CDR-L2, and CDR-L3 on the light chain (Presta, 1992; Al-Lazikani et al., 1997).
Our work represents each single protein residue in terms of the residue type $s_{i}\in\left\{ACDEFGHIKLMNPQRSTVWY\right\}$, the coordinate ${x}_{i}\in\mathbb{R}^{3}$, and the orientation $\textbf{O}_{i}\in\text{SO}\left(3\right)$, where $i=1,...,N$ and N is the number of residues in the complex. Concretely, assuming the CDR sequence to be generated includes $m$ amino acids and starts from position $a$, it can be denoted as $R=\left\{s_{j}\mid j\in\{a+1,...,a+m\}\right\}$. Let $M$ be the length of the antibody, the antibody framework is defined as $C_{ab}=\left\{\left(s_{i},x_{j},\textbf{O}_{j}\right)\mid i\in\left\{1,...,M\right\}\backslash\{a+1,...,a+m\},j\in\left\{1,...,M\right\}\right\}$. The antibody framework sequence is defined as $S_{ab}=\left\{s_{i}\mid i\in\left\{1,...,M\right\}\backslash\{a+1,...,a+m\}\right\}$. The antigen is defined as $C_{ag}=\left\{\left(s_{i},x_{i},\textbf{O}_{i}\right)\mid i\in\left\{M+1,...,N\right\}\right\}$. The retrieved CDR-like fragments are defined as ${\mathbb{A}}=\left\{A_{i}\mid i\in\left\{1,...,k\right\}\right\}$. The goal of our framework is to extract the antibody CDR structure from the antibody framework complex $C_{ab}$, then input it into the retrieval module to retrieve ${\mathbb{A}}$, and ultimately predict the distribution of $R$ through $C_{ag}$, $C_{ab}$ and ${\mathbb{A}}$.
@@ 3.2 Diffusion model for antibody design
Due to diffusion models’ excellent performance and controllability, there are now many diffusion-based works that have achieved notable results (Luo et al., 2022; Villegas-Morcillo et al., 2023; Kulytė et al., 2024). To be specific, they are denoising probabilistic diffusion models that transform the amino acid type $s$, the backbone $C\alpha$ atom coordinates $x$, and the amino acid orientation O during the diffusion process. We focus on sequence, of which the forward process perturbs the data in the following ways (Hoogeboom et al., 2021):
@@
To reverse the aforementioned forward process and denoise to generate CDR sequence, predictions need to be made by a neural network $F$(·)$[j]$, which takes the antibody-antigen context as condition:
@@
As an example, this work uses Diffab (Luo et al., 2022) as the backbone for the generative model to conduct retrieval augmented generation. Note that the proposed retrieval system is generative model agnostic, and the developed modules can be integrated with any diffusion generative model.
@@ 4 Methods
Figure 2: The overall architecture of the proposed RADAb framework. (A) Structural retrieval process, the CDR backbone is input into MASTER and the output is a set of ranked CDR-like fragments. (B) Diffusion process and denoising network which takes antibody-antigen context and retrieved evolutionary information as conditions. (C) Our method restricts the antibody to a small region through fixed structural constraints and retrieval-augmented constraints (functional constraints) to achieve higher fitness.
@@ 4.1 Structural retrieval of CDR fragments
To balance the quality of results and the retrieval speed, we use MASTER (Zhou & Grigoryan, 2015) for the search. MASTER uses the root-mean-square deviation (RMSD) of backbone atoms as a similarity measure. It queries structural fragments composed of one or more non-contiguous segments and can find all matching fragments from the database within a given RMSD threshold. This allows for fast and accurate searches in the PDB database for protein motifs. Note that MASTER can utilize only the backbone information without any leakage of sequence data during the search process. The retrieval procedure is described in Algorithm 1 and detailed in Appendix A.3.
@@
18:         insert match $A$ into ${\mathbb{A}}$
@@ 4.2.1 Global geometry context information branch
Evolutionary encoder Recent advances have shown the structure-informed protein language model (PLM) is an excellent tool for creating protein sequence embeddings and providing evolutionary information (Zheng et al., 2023; Shanker et al., 2024). Thus, we take ESM2 (Lin et al., 2023) as an antibody sequence encoder, aiming to capture the evolutionary relations of antibody residues. The state of antibody sequence with CDR at timestep $t$ is fed into it and output is defined as $e^{t}$.
Structure-informed network The above encoding is used as conditional input to the Structure-informed network. They, along with the CDR sequence and structural state at the current time step, will be input into a stack of Invariant Point Attention (IPA) (Jumper et al., 2021) layers, and jointly transform into a hidden representation $h_{i}$. Subsequently, the hidden representation $h_{i}$ is transformed by an MLP to obtain the probability representation $r_{\text{global}}$ of the amino acid type at each CDR site. This probability representation is then input to the local CDR-focused branch.
@@ 4.2.2 Local CDR-focused information branch
CDR-focused Axial Attention The local CDR-like branch is constituted of a stack of axial attention layers, referred to as CDR-focused Axial Attention. Given that the CDR-like fragments exhibit structures similar to actual CDRs, we employed a tied row attention mechanism used in MSATransformer (Rao et al., 2021) to leverage these retrieval results. In the standard axial attention (Ho et al., 2019) mechanism, each row and column are considered independently. However, in MSA (Multiple Sequence Alignment), each sequence exhibits relatively similar structural features. Our matrix format is well-suited for adopting a tied row attention mechanism to fully utilize the structural similarity. When calculating the attention scores for each row, this mechanism simultaneously considers the scores of other rows. This approach not only leverages the structural similarity but also reduces memory usage.
@@
Skip connection for information fusion Although the probability distribution of the CDR region created by the antigen-antibody context features has already been fed into the network, to prevent the loss of antigen-antibody context information during forward propagation, the embedding $r_{\text{local}}$ and $r_{\text{global}}$ are added by a skip connection module (He et al., 2016), then execute softmax to obtain the final probability distribution.
@@ 4.3 Model training and inference
[…]
Conditional reverse diffusion process We employ DDPM to generate sequences. The model starts from time step T, initializing each site of the antibody CDR region as a uniform distribution. Then, through the frozen ESM encoder $E$(·), learned global context network $F$(·)$[j]$ and the local CDR-focused network $G$(·)$[j]$, they predict the noise distribution at each time step jointly and denoise step-by-step:
@@ 5 Experiments
The dataset for training the model is obtained from the SAbDab and our established CDR-like fragments dataset. Following the previous work (Luo et al., 2022), we first eliminated structures with a resolution lower than 4Å and removed antibodies that target non-protein antigens. Chothia (Chothia & Lesk, 1987) in ANARCI (Dunbar & Deane, 2016) is used for renumbering antibody residues. We clustered the SADab datasets based on 50% sequence similarity in the CDR-H3 region, and chose 50 PDB files comprising 63 antibody-antigen complex structures as the test set. To ensure distinct training and test sets, we removed structures from the training set that were part of the same clusters as those in the test set.

### v2

@@ 1 Introduction
Antibodies, essential Y-shaped proteins in the immune system, are pivotal for recognizing and neutralizing specific pathogens known as antigens. This specificity primarily arises from the Complementarity Determining Regions (CDRs), which are crucial for binding affinity to antigens (jones1986replacing; ewert2004stability; xu2000diversity; akbar2021compact). The design of effective CDRs is therefore central to developing potent therapeutic antibodies, a dominant class of protein therapeutics. However, the development of these antibodies typically relies on labor-intensive experimental methods such as animal immunization or screening extensive antibody libraries, often failing to produce antibodies that target therapeutically relevant epitopes effectively. Thus, the ability to generate new antibodies with pre-defined biochemical properties in silico carries the promise of speeding up the drug design process.
Computational efforts in antibody design have traditionally involved grafting residues onto existing structures (sormanni2015rational), sampling alternative native CDR loops to enhance affinities(aguilar2022fragment), and using tools like Rosetta for sequence design improvements in interacting regions (adolf2018rosettaantibodydesign). Many recent studies have focused on applying deep generative models to design antibodies (luo2022antigen; martinkus2024abdiffuser; zhuabx). They take advantage of geometric learning and generative models to capture the higher-order interactions among residues directly from the data. These innovations provide more efficient methods to search sequence and structure spaces.
Albeit powerful, current generative models struggle to design antibodies that adhere to structural constraints and exhibit desired biological properties. This challenge primarily arises from a lack of diversity in the available training data. Predominantly, research efforts have relied on the SAbDab database (dunbar2014sabdab), which comprises fewer than ten thousand antigen-antibody complex structures. The limited scope of this dataset restricts the models’ ability to capture comprehensive high-order interaction information between antigen-antibody residues, thereby increasing the risk of overfitting. Moreover, most existing methodologies attempt to design antibody sequences de novo, without the benefit of template-based guidance. This approach inherently demands a greater volume of data and extensive training or fine-tuning on specific datasets to achieve efficacy in practical applications.
@@
To this end, we introduce the Retrieval-Augmented Diffusion Antibody design model (RADAb), a novel semi-parametric antibody design framework. To fully exploit the protein structure space, we first compiled a database of CDR-like fragments from the non-redundant Protein Data Bank (PDB) (berman2000protein). These CDR-like fragments are linear functional motifs structurally compatible with an antibody CDR loop, found in any protein geometry within the PDB. For a given antibody to be improved, we perform a structural retrieval to obtain motifs with structures similar to the desired CDR framework. As protein sequences capable of folding into similar structures often share homology and consensus, we hypothesize that these retrieved motifs, enriched with evolutionary information, can enhance the model’s generalization.
[…]
@@ 2 Related work
Antibody Design Computational antibody design primarily follows two paths: conventional energy function optimization methods and machine learning approaches. Early antibody design methods were often limited to sequence similarity and energy function optimization (lapidoth2015abdesign; adolf2018rosettaantibodydesign). Recent success of machine learning approaches mainly falls into two directions: antibody sequence design and antigen-specific antibody sequence-structure co-design. The methods used for antibody sequence design mainly include language-based models (ruffolo2021deciphering; olsen2022ablang; wang2023pre) and inverse folding models (dreyer2023inverse; hoie2024antifold). The other line focuses on antibody sequence-structure co-design mainly taking antibody-antigen complex as a graph, then using graph networks to extract features and predict the coordinates and residue type of antibody CDR (jiniterative2021; kongMEAN; kong2023end; lin2024geoab; luo2022antigen; zhuabx; martinkus2024abdiffuser). While these works are undoubtedly powerful, they often generate antibodies from scratch without incorporating explicit structure constraints, which can introduce challenges in designing functional antibodies (zhou2024antigen). Instead, our method leverages the power of templates from a structure-informed perspective.
@@
Diffusion generative models Diffusion models (sohl2015deep; song2020denoising; ho2020denoising) are a class of generative models that have achieved impressive progress on a lot of generation tasks. Denoising diffusion probabilistic models (DDPMs) are a branch of diffusion models, which contain two Markov processes. The forward process perturbs the data into pure noise, and then learns to generate data by reversing the forward Markov process. Because of the diffusion model’s flexibility and controllability, numerous works are focusing on employing retrieval-augmented methods to complement the diffusion framework for text-to-image generation (sheyninknn), image generation (blattmann2022retrieval), human motion generation (zhang2023remodiffuse) and small molecule generation (huanginteraction).
@@
Retrieval augmented generative models Retrieval augmented generation technique was first proposed in the field of natural language processing to enhance the language models by introducing an additional database (lewis2020retrieval; pmlr-v119-guu20a), prompting the language models to generate more realistic and diverse results. Subsequently, retrieval augmented generation (RAG) has conducted diverse explorations in large fields, including natural language processing (zhang2023retrieve; gao2023retrieval; xuretrieval; yoranmaking; caffagni2024wiki) and computer vision (long2022retrieval; blattmann2022retrieval; hu2023reveal; rao2023retrieval).
Recently, several studies have been proposed using retrieval techniques to enhance molecular generation. RetMol (wang2023retrievalbased) generates new molecules based on existing small molecules by retrieving a set of exemplar moleculars. IRDiff (huanginteraction) enhances protein-specific molecular generation by using protein pockets to retrieve moleculars which interct with the pocket. Although there has been some retrieval-based work in the field of protein design and discovery (zhou2015rapid; aguilar2022fragment), to the best of our knowledge, this is the first-of-its-kind retrieval-based generative framework for antibody design.
@@ 3.1 Notations
Antibody consists of two heavy chains and two light chains. Each chain’s tip has a complementary site that specifically binds to a unique epitope on the antigen. This site includes six complementary-determining regions (CDRs): CDR-H1, CDR-H2, and CDR-H3 on the heavy chain, and CDR-L1, CDR-L2, and CDR-L3 on the light chain (presta1992antibody; al1997standard).
Our work represents each single protein residue in terms of the residue type $s_{i}\in\left\{ACDEFGHIKLMNPQRSTVWY\right\}$, the coordinate ${x}_{i}\in\mathbb{R}^{3}$, and the orientation $\textbf{O}_{i}\in\text{SO}\left(3\right)$, where i=1,,Ni=1,...,N and N is the number of residues in the complex. Concretely, assuming the CDR sequence to be generated includes $m$ amino acids and starts from position $a$, it can be denoted as $R=\left\{s_{j}\mid j\in\{a+1,...,a+m\}\right\}$. Let $M$ be the length of the antibody, the antibody framework is defined as $C_{ab}=\left\{\left(s_{i},x_{j},\textbf{O}_{j}\right)\mid i\in\left\{1,...,M\right\}\backslash\{a+1,...,a+m\},j\in\left\{1,...,M\right\}\right\}$. The antibody framework sequence is defined as $S_{ab}=\left\{s_{i}\mid i\in\left\{1,...,M\right\}\backslash\{a+1,...,a+m\}\right\}$. The antigen is defined as $C_{ag}=\left\{\left(s_{i},x_{i},\textbf{O}_{i}\right)\mid i\in\left\{M+1,...,N\right\}\right\}$. The retrieved CDR-like fragments are defined as ${\mathbb{A}}=\left\{A_{i}\mid i\in\left\{1,...,k\right\}\right\}$. The goal of our framework is to extract the antibody CDR structure from the antibody framework complex $C_{ab}$, then input it into the retrieval module to retrieve $\mathbb{A}$}, and ultimately predict the distribution of $R$ through $C_{ag}$, $C_{ab}$ and $\mathbb{A}$}.
@@ 3.2 Diffusion model for antibody design
Due to diffusion models’ excellent performance and controllability, there are now many diffusion-based works that have achieved notable results (luo2022antigen; villegas-morcillo2023guiding; kulytė2024improving). To be specific, they are denoising probabilistic diffusion models that transform the amino acid type $s$, the backbone $C\alpha$ atom coordinates $x$, and the amino acid orientation O during the diffusion process. We focus on sequence, of which the forward process perturbs the data in the following ways (multHoo):
@@
To reverse the aforementioned forward process and denoise to generate CDR sequence, predictions need to be made by a neural network FF(·)$[j]$, which takes the antibody-antigen context as condition:
@@
As an example, this work uses Diffab (luo2022antigen) as the backbone for the generative model to conduct retrieval augmented generation. Note that the proposed retrieval system is generative model agnostic, and the developed modules can be integrated with any diffusion generative model.
@@ 4 Methods
Figure 2: The overall architecture of the proposed RADAb framework. (A) Structural retrieval process, the CDR backbone is input into MASTER and the output is a set of ranked CDR-like fragments. (B) Diffusion process and denoising network which takes antibody-antigen context and retrieved evolutionary information as conditions. The structure is fixed during diffusion process. (C) Our method restricts the antibody to a small region through fixed structural constraints and retrieval-augmented constraints (functional constraints) to achieve higher fitness.
@@ 4.1 Structural retrieval of CDR fragments
To balance the quality of results and the retrieval speed, we use MASTER (zhou2015rapid) for the search. MASTER uses the root-mean-square deviation (RMSD) of backbone atoms as a similarity measure. It queries structural fragments composed of one or more non-contiguous segments and can find all matching fragments from the database within a given RMSD threshold. This allows for fast and accurate searches in the PDB database for protein motifs. Note that MASTER can utilize only the backbone information without any leakage of sequence data during the search process. The retrieval procedure is described in Algorithm 1 and detailed in Appendix A.3.
@@
18:   insert match AA into $\mathbb{A}$}
@@ 4.2.1 Global geometry context information branch
Evolutionary encoder Recent advances have shown the structure-informed protein language model (PLM) is an excellent tool for creating protein sequence embeddings and providing evolutionary information (zheng2023structure; shanker2024unsupervised). Thus, we take ESM2 (lin2023evolutionary) as an antibody sequence encoder, aiming to capture the evolutionary relations of antibody residues. The state of antibody sequence with CDR at timestep $t$ is fed into it and output is defined as $e^{t}$.
Structure-informed network The above encoding is used as conditional input to the Structure-informed network. They, along with the CDR sequence and structural state at the current time step, will be input into a stack of Invariant Point Attention (IPA) (jumper2021highly) layers, and jointly transform into a hidden representation $h_{i}$. Subsequently, the hidden representation $h_{i}$ is transformed by an MLP to obtain the probability representation $r_{\text{global}}$ of the amino acid type at each CDR site. This probability representation is then input to the local CDR-focused branch.
@@ 4.2.2 Local CDR-focused information branch
CDR-focused Axial Attention The local CDR-like branch is constituted of a stack of axial attention layers, referred to as CDR-focused Axial Attention. Given that the CDR-like fragments exhibit structures similar to actual CDRs, we employed a tied row attention mechanism used in MSATransformer (rao2021msa) to leverage these retrieval results. In the standard axial attention (ho2019axial) mechanism, each row and column are considered independently. However, in MSA (Multiple Sequence Alignment), each sequence exhibits relatively similar structural features. Our matrix format is well-suited for adopting a tied row attention mechanism to fully utilize the structural similarity. When calculating the attention scores for each row, this mechanism simultaneously considers the scores of other rows. This approach not only leverages the structural similarity but also reduces memory usage.
@@
Skip connection for information fusion Although the probability distribution of the CDR region created by the antigen-antibody context features has already been fed into the network, to prevent the loss of antigen-antibody context information during forward propagation, the embedding $r_{\text{local}}$ and $r_{\text{global}}$ are added by a skip connection module (he2016deep), then execute softmax to obtain the final probability distribution.
@@ 4.3 Model training and inference
$L=E_{t\sim\,\,\text{Uniform\,\,}\left(1...T\right)}L_{\text{type\,\,}}^{t}$ (6)
Conditional reverse diffusion process We employ DDPM to generate sequences. The model starts from time step T, initializing each site of the antibody CDR region as a uniform distribution. Then, through the frozen ESM encoder EE(·), learned global context network FF(·)$[j]$ and the local CDR-focused network GG(·)$[j]$, they predict the noise distribution at each time step jointly and denoise step-by-step:
@@ 5 Experiments
The dataset for training the model is obtained from the SAbDab and our established CDR-like fragments dataset. Following the previous work (luo2022antigen), we first eliminated structures with a resolution lower than 4Å and removed antibodies that target non-protein antigens. Chothia (chothia1987canonical) in ANARCI (dunbar2016anarci) is used for renumbering antibody residues. We clustered the SADab datasets based on 50% sequence similarity in the CDR-H3 region, and chose 50 PDB files comprising 63 antibody-antigen complex structures as the test set. To ensure distinct training and test sets, we removed structures from the training set that were part of the same clusters as those in the test set.